If (x) =4x+3,Which of the following is true and why:
1.f(x+h)=4x+3+h
2.f(x+h)=4x+4h+3
3.f(x+h)4x+4h+6
can you tell me I'll appreciate thank you for this question and the previous one

Answer by stanbon(75887)

If (x) =4x+3,Which of the following is true and why:
Substitute "x+h" in place of "x" to get:
f(x+h) = 4(x+h)+3 = 4x+4h+3
----------------------------------
1.f(x+h)=4x+3+h
2.f(x+h)=4x+4h+3
3.f(x+h)4x+4h+6
---------------------
Ans: 2
Cheers,
Stan H.